# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v1

@@ C.1. Quantitative Results
We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 9 and 10. Experiments show that the proposed CTAOD can largely maintain strong performance in the long term.
@@ C.3. Additional Ablation Study
Table 7. Additional Ablation Study. The “FT” represents replacing dynamic threshold in CTAOD with a fixed threshold. The “SR” denotes the stochastic restoration (Wang et al., 2022), and the “DR” indicates the data-driven restoration (Brahma and Rai, 2023). All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
CTAOD FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean
@@
As presented in Table 7, we perform an additional ablation study to evaluate the effectiveness of the proposed Dynamic Threshold (DT) and Data-driven Stochastic Restoration (DSR). The findings indicate that the DT approach results in improvements of 1.3, 0.8, and 0.5 mAP over fixed thresholds set at 0.9, 0.8, and 0.7, respectively. While the stochastic restoration and data-driven restoration limit the performance to 27.1 and 28.3 mAP, the DSR mechanism raises the performance to 29.0 mAP.
Appendix D Limitation
Table 8. Detailed hyper-parameters for each benchmark. “City2City-C-Short” denotes the short-term Cityscapes-to-Cityscapes-C task, “City2City-C-Long” denotes long-term Cityscapes-to-Cityscapes-C task, and “ City2ACDC-Long” denotes long-term Cityscapes-to-ACDC task.
@@ Appendix D Limitation
$\beta$ Hyper-parameter in DT 0.95 0.95 0.95 0.95
@@
$q$ Quantile value 0.05 0.001 0.001 0.01
@@
1:Initialization: Object detectors: The student and teacher model parameters $\theta_{0}^{std}$ and $\theta_{0}^{tch}$ initialized from source-trained model $\theta_{0}$. Hyper-parameters: learning rates $\gamma$, Momentum $\alpha$ in exponential moving average (EMA), temperature $\tau$, thresholds $\delta_{0}^{c}$, $\beta$ and $\epsilon$ in dynamic threshold, and quantile value $q$.
@@
6:       // 1. Object-level Contrastive Learning
7:       Extract teacher features $F_{t}^{T}$ and student features $F_{t}^{S}$ based on $p_{t}$ using RoIAlign.
8:       Compute contrastive learning loss $\mathcal{L}_{cl}$:
9:             $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)}$
10:       // 2. Dynamic Threshold
11:       for each category $c$ do
12:             Compute the mean predicted scores $\overline{l_{t}^{c}}$ of category $c$
13:             Update dynamic thresholds based on predictions:
14:                   $\delta_{t}^{c}\leftarrow\beta\cdot\delta_{t-1}^{c}+(1-\beta)\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}}$
15:       end for
16:       Generate the pseudo-label $\hat{y}_{t}$ through the category-specific threshold $\delta_{0}^{c}$
17:       Compute the supervised loss $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t})$
18:       Perform gradient descent to update the student model $\theta_{t}^{std}$ and get the gradient matrix $G_{t}$
19:       Update teacher model $\theta_{t}^{tch}$ via EMA strategy
20:       // 3. Data-driven Stochastic Restoration
21:       Generate random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$
22:       Generate the fisher information matrix: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$
23:       Find the $q$-quantile of $F_{t}$: $\eta=quantile(F_{t},q)$
24:       Generate the mask matrix: $M_{t}=F_{t}<\eta$
25:       Reset the updated student model: $\theta_{t}^{std}=M_{t}\odot\theta_{0}^{std}+(1-M_{t})\odot\theta_{t}^{std}$
26:end for
27:Output: Teacher predictions $y_{t}$
Time $t\xrightarrow{\hskip 403.26341pt}$
Round 1 2 3 4 5 /
Tent 35.8 8.1 0.2 29.2 6.2 34.6 7.7 0.1 27.5 5.5 33.5 7.8 0.1 25.5 5.1 32 7.3 0.1 23.2 4.6 30.5 6.8 0.1 21.0 3.9 / /
Ours 39.0 10.4 0.8 48.0 13.8 48.2 13.7 3.1 52.9 18.1 49.3 15.9 5.3 53.3 21.3 48.5 17.1 7.1 52.2 23.5 49.4 18.6 7.7 51.7 25.0 / /
Round 6 7 8 9 10 Mean Gain
Tent 28.6 6.4 0.1 19.0 3.5 26.6 6.3 0.1 16.9 3.0 24.8 5.9 0.1 15.2 2.7 22.9 5.6 0.1 13.7 2.7 20.4 5.0 0.1 11.5 2.4 12.0 -4.4
Ours 48.5 19.7 8.9 50.7 26.3 47.8 19.6 9.8 49.0 26.1 45.7 19.3 10.5 49.4 26.9 45.2 19.7 11.2 47.4 26.9 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
Time $t\xrightarrow{\hskip 390.25534pt}$
Round 1 2 3 4 5 /
Tent 52.4 18.6 33.4 38.9 52.4 18.1 32.5 37.7 52.2 17.8 32.7 36.7 51.7 17.4 31.4 36 50.4 16.5 31.7 32.9 / /
Ours 52.9 20.4 34.5 42.9 53 22.9 38.6 44.6 54.2 24 39.5 45.3 54.5 23.8 39.6 45.8 54.4 24.4 39.2 45.4 / /
Round 6 7 8 9 10 Mean Gain
Tent 49 15.8 29.5 30.6 45.8 14.6 28.1 28.5 42.4 12.6 25.8 24.9 38.9 11 23.6 21.6 35 9.5 21.5 18.3 30.5 -5.5
Ours 53.9 24.9 39.2 45.1 53.5 25.1 39 45.3 53.3 24.6 39.3 45.3 53.2 24.5 39.3 44.9 52.5 24.3 38.8 44.7 40.3 +4.3

### v3

@@ C.1. Quantitative Results
We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 8 and 9. Experiments show that the proposed CTAOD can largely maintain strong performance in the long term.
@@ C.3. Additional Ablation Study
CTAOD-unstop FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean
@@
Table 6. Additional Ablation Study. The “CTAOD-unstop” denotes CTAOD without dynamic skipping. The “FT” represents replacing dynamic threshold in CTAOD with a fixed threshold. The “SR” denotes the stochastic restoration (Wang et al., 2022), and the “DR” indicates the data-driven restoration (Brahma and Rai, 2023). All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
As presented in Table 6, we perform an additional ablation study to evaluate the effectiveness of the proposed Dynamic Threshold (DT) strategy in Adaptive Monitoring and Data-driven Stochastic Restoration (DSR). The findings indicate that the DT approach results in improvements of 1.3, 0.8, and 0.5 mAP over fixed thresholds set at 0.9, 0.8, and 0.7, respectively. While the stochastic restoration and data-driven restoration limit the performance to 27.1 and 28.3 mAP, the DSR mechanism raises the performance to 29.0 mAP.
@@
$\beta_{t}$ Hyper-parameter in DT 0.95 0.95 0.95 0.95
@@
$\beta_{s}$ Hyper-parameter in DS 0.7 0.75 0.6 0.7
$\delta_{s}$ The threshold in DS 1.5 1.4 1.4 1.4
$q$ Quantile value 0.01 0.001 0.001 0.01
Table 7. Detailed hyper-parameters for each benchmark. “DT” and “DS” denotes dynamic threshold and dynamic skipping in adaptive monitoring. “City2City-C-Short” denotes the short-term Cityscapes-to-Cityscapes-C task, “City2City-C-Long” denotes long-term Cityscapes-to-Cityscapes-C task, and “ City2ACDC-Long” denotes long-term Cityscapes-to-ACDC task.
Time $t\xrightarrow{\hskip 403.26341pt}$
Round 1 2 3 4 5 /
Ours-unstop 39.0 10.4 0.8 48.0 13.8 48.2 13.7 3.1 52.9 18.1 49.3 15.9 5.3 53.3 21.3 48.5 17.1 7.1 52.2 23.5 49.4 18.6 7.7 51.7 25.0 / /
Ours 39.2 9.9 0.8 46.7 12.7 46.8 13.5 2.7 52.3 17.9 49.1 15.8 4.5 53.9 21.2 48.8 17.7 6.9 53.8 23.1 49.4 18.5 8.0 52.9 24.7 / /
Round 6 7 8 9 10 Mean Gain
Ours-unstop 48.5 19.7 8.9 50.7 26.3 47.8 19.6 9.8 49.0 26.1 45.7 19.3 10.5 49.4 26.9 45.2 19.7 11.2 47.4 26.9 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
Ours 48.6 19.8 9.1 51.2 25.4 48.8 20.1 10.0 49.5 27.0 47.3 19.9 10.0 49.2 27.9 46.9 19.4 11.0 47.5 28.9 46.2 19.5 11.3 47.3 29.1 29.2 +12.8
Appendix D Additional Related Work
D.1. Continual Learning
Appendix E Limitation
Time $t\xrightarrow{\hskip 390.25534pt}$
Round 1 2 3 4 5 /
Ours-unstop 52.9 20.4 34.5 42.9 53 22.9 38.6 44.6 54.2 24 39.5 45.3 54.5 23.8 39.6 45.8 54.4 24.4 39.2 45.4 / /
Ours 52.8 20.1 35.2 42.9 53.6 22.2 38.3 44.2 53.9 23.5 39.6 45.0 54.4 24.6 38.6 45.5 54.6 25.1 39.2 46.3 / /
Round 6 7 8 9 10 Mean Gain
Ours-unstop 53.9 24.9 39.2 45.1 53.5 25.1 39 45.3 53.3 24.6 39.3 45.3 53.2 24.5 39.3 44.9 52.5 24.3 38.8 44.7 40.3 +4.3
Ours 54.4 25.4 38.7 46.1 53.5 25.6 38.5 46.2 54.1 25.9 38.0 45.8 53.3 25.9 37.9 45.4 53.5 25.4 37.4 45.3 40.4 +4.4
@@ Appendix E Limitation
1:Initialization: Object detectors: The student and teacher model parameters $\theta_{0}^{std}$ and $\theta_{0}^{tch}$ initialized from source-trained model $\theta_{0}$. Hyper-parameters: learning rates $\gamma$, Momentum $\alpha$ in exponential moving average (EMA), temperature $\tau$, thresholds $\delta_{0}^{c}$, $\beta_{t}$ and $\epsilon$ in dynamic threshold, $\beta_{s}$ and $\delta_{s}$ in dynamic skipping, and quantile value $q$.
@@
6:       // 1. Adaptive Monitoring
7:       Compute the mean predicted scores of all instances $\overline{l}_{t}$ and the index $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$
8:       Udapate the exponentially moving average of mean scores $\overline{l}_{t}^{ema}$ through:
9:             $\overline{l}_{t}^{ema}\leftarrow\beta_{s}\cdot\overline{l}_{t-1}^{ema}+(1-\beta_{s})\cdot\overline{l}_{t},$
10:       if $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$ not lies with $(\frac{1}{\delta_{s}},\delta_{s})$ then
11:             Pause the current adaption
12:       end if
13:       for each category $c$ do
14:             Compute the mean predicted scores $\overline{l_{t}^{c}}$ of category $c$
15:             Update dynamic thresholds based on predictions:
16:                   $\delta_{t}^{c}\leftarrow\beta_{t}\cdot\delta_{t-1}^{c}+(1-\beta_{t})\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}}$
17:       end for
18:       // 2. Object-level Contrastive Learning
19:       Extract teacher features $F_{t}^{T}$ and student features $F_{t}^{S}$ based on $p_{t}$ using RoIAlign
20:       Compute contrastive learning loss $\mathcal{L}_{cl}$:
21:             $\mathcal{L}_{cl}(x_{t})=\frac{1}{l}\sum_{i=1}^{l}-\log\frac{\exp(f_{i}^{T}\cdot f_{i}^{S}/\tau)}{\sum_{j=1}^{l}\exp(f_{i}^{T}\cdot f_{j}^{S}/\tau)}$
22:       Generate the pseudo-label $\hat{y}_{t}$ through the category-specific threshold $\delta_{0}^{c}$
23:       Compute the supervised loss $\mathcal{L}_{pl}(x_{t})=\mathcal{L}_{rpn}(x_{t},\hat{y}_{t})+\mathcal{L}_{rcnn}(x_{t},\hat{y}_{t})$
24:       Perform gradient descent to update the student model $\theta_{t}^{std}$ and get the gradient matrix $G_{t}$
25:       Update teacher model $\theta_{t}^{tch}$ via EMA strategy
26:       // 3. Data-driven Stochastic Restoration
27:       Generate random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$
28:       Generate the fisher information matrix: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$
29:       Find the $q$-quantile of $F_{t}$: $\eta=quantile(F_{t},q)$
30:       Generate the mask matrix: $M_{t}=F_{t}<\eta$
31:       Reset the updated student model: $\theta_{t}^{std}=M_{t}\odot\theta_{0}^{std}+(1-M_{t})\odot\theta_{t}^{std}$
32:end for
33:Output: Teacher predictions $y_{t}$